OpenStudy (anonymous):

please write expression in a+bi? 6+2i/9-3i

OpenStudy (anonymous):

To simplify it multiply top and bottom by the conjugate of the bottom ... so by 9 + 3i: = [(6 + 2i)(9 + 3i)] / [(9 - 3i)(9 + 3i)] = [54 + 18i + 18i + 6i²] / [9² - 9i²] = [54 + 36i - 6] / [81 + 9] ... [b/c i² = -1] = [48 + 36i] / 90 = (48/90) + (36i / 90) = (8/15) + (6/15)i = (8/15) + (2/5)i

OpenStudy (anonymous):

conjugate formula?

OpenStudy (anonymous):